Eidetic Variations on Viewing and Understanding

Greg Hudson

November 18, 1994

“Viewing” and “understanding” are two similar phenomenological terms which we have approached from different angles: viewing is the ontological phenomenon of which staring at, remembering, and considering are ontic modes; understanding is a rational ontic mode of disclosure, which is a process of becoming familiar with a situation.

The essential difference between viewing and understanding is that viewing implies only a direction of attention upon a situation, while understanding involves a change in level of familiarity with a situation. To illustrate this, consider the act of staring at a tape recorder. If I do not already understand tape recorders, the most I can do by staring at the tape recorder is size it up, and beyond that, I will not develop any further understanding of it. Yet, I am still viewing the tape recorder by staring at it.

Thus we can have viewing without understnding. It is a bit more far-fetched to consider understanding without viewing; we might try to familiarize ourselves with a situation by examining the circumstances surrounding it without viewing the situation itself, but this probably qualifies as another ontic mode of viewing. For instance, when we test a piece of software by treating it as a black box, we try not to consider the implementation of the software, but we are probably still viewing it in our attempt to gain understanding of it.